OpenStudy (anonymous):

ahhh please help! (x − 3)^2

OpenStudy (hba):

Use the formula (a-b)^2=a^2-2ab+b^2 a=x b=3 best of luck :)

OpenStudy (anonymous):

im still a little confused can you work through it with me

OpenStudy (hba):

Yes you start working and i will help you where you go wrong :)

OpenStudy (anonymous):

can you draww the formla please

OpenStudy (hba):

@agiles The formula is up there you just have to use it :)

OpenStudy (hba):

@agiles I am getting some news from users that you are messaging for help ?

OpenStudy (hba):

@agiles Comeone do some work !!

OpenStudy (uri):

HAAAAAAAY,WELCOME TO OPENSTUDY @agiles :D

OpenStudy (anonymous):

Just use FOIL (First Outer Inner Last) basically (x-3)(x-3)

OpenStudy (anonymous):

Thank you so much @MissMathmatics